OpenStudy (3psilon):

What is the additive inverse and multiplicative inverse of 3-i ? Can someone explain to me what those are please?

OpenStudy (helder_edwin):

the additive inverse of x is -x. the multiplicative inverse of \(x\neq0\) is 1/x

OpenStudy (helder_edwin):

so for instance. if \(z=-3+4i\), then \(-z=3-4i\) is the additive inverse of z. since \(z=-3+4i\neq0\) then its multiplicative inverse is \[ \large z^{-1}=\frac{1}{-3+4i}\cdot\frac{-3-4i}{-3-4i}=\frac{-3-4i}{9+16}= \frac{-3-4i}{25}=-\frac{3}{25}-\frac{4}{25}i \]

OpenStudy (3psilon):

why do we multiply by the conjugate again @helder_edwin ?

OpenStudy (3psilon):

I get everything else I just forgot that concept

OpenStudy (helder_edwin):

it is a "shortcut" to do division with complex numbers.

OpenStudy (3psilon):

Ohh okay okay thank you so much!